OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) 2x − 6 i already know that 2(x-3) =2x-6 =6/2 = x=3

OpenStudy (science0229):

Using the distributive property, which states that a(b+c)=ab+ac, 2*x is 2x and 2*-3 is -6, so the left is equal to the right.

OpenStudy (anonymous):

why ?

OpenStudy (science0229):

There is a real number property that states "For any real numbers a,b, and c, the following is always true: a(b+c)=ab+ac", and it's called the Distributive Property.

OpenStudy (science0229):